# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v4

@@ VI-D4 Evaluation Challenges
For training-free token compression methods, current evaluations often rely on metrics such as the number of compressed tokens and FLOPs. However, these metrics offer an incomplete picture. While the number of compressed tokens provides a preliminary classification, the location where compression is applied significantly impacts the downstream computational load; earlier pruning generally leads to greater reductions. Similarly, FLOPs, while useful for theoretical computational estimates, frequently do not accurately reflect actual inference speed. Therefore, for training-free methods, more practical metrics like Time To First Token (TTFT) and decoding latency per token are crucial for a more accurate assessment of real-world inference acceleration.
Evaluation Benchmarks Gap. Current evaluation datasets for MLLM token compression, often relying on general multimodal benchmarks, provide insufficient granularity. For example, in challenging long video understanding tasks, performance hinges more on sparse frame sampling, capturing key frames, than on the specific token compression method. This can obscure the true impact of token compression, making its efficacy appear negligible.
This reveals a critical gap: current datasets often fail to isolate and precisely measure the effect of token compression. Therefore, adopting evaluation methodologies like EffiVLM-Bench [198], which focuses on training-free acceleration evaluation, is crucial for accurately assessing the true efficacy and nuanced benefits of token compression methods.
VI-E Pruning Location and Trade-offs
@@
VI-F Future Directions
VI-F1 Unified Multimodal Token Compression
Although different modalities exhibit unique redundancy patterns requiring specialized compression strategies, contemporary MLLMs predominantly focus on joint multimodal inference [10], whereas unimodal scenarios remain constrained by single-modality inputs. As established in Sections III, IV, and V, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. Furthermore, leveraging cross-modal correlations represents a promising research direction for the future.
VI-F2 Improved Architecture
Current token compression methods are often employed as a remedial measure to efficiently process long contexts. However, a more valuable approach might involve designing model architectures that intrinsically account for data redundancy during their initial conception. By doing so, the number of tokens could be reduced during the abstraction of data features. This is particularly relevant for current architectures, especially those of video LLMs, where generated tokens still exhibit significant redundancy. Therefore, exploring architectural designs that inherently foster more condensed information abstraction from the outset represents a promising research direction.
VII Applications
@@
VII-A GUI Agents and Human-Computer Interaction
@@
Graphical user interface (GUI) agents perceive and interact with visual interfaces, interpret natural language instructions, analyze GUI states, and execute corresponding actions. These agents have to parse screen streams in real-time, producing extensive token sequences that often exceed computational limits [199, 148]. Multimodal token compression enhances the efficiency of GUI agents. This approach mitigates context overflow in extended operation sequences by dynamically compressing redundant visual elements (e.g., extra white space or simple backgrounds). For some small but important control elements, it should also eliminate other irrelevant visual elements and highlight their importance. For instance, ShowUI [41] is the first model to apply token selection strategy to GUI agents. ShowUI segments GUI screenshots into connected components by clustering pixels with similar RGB values, significantly reducing the total number of discrete elements. During both training and inference phases, the system employs an adaptive token selection strategy that probabilistically prunes redundant tokens within these components, thereby optimizing computational efficiency while preserving functional semantics However, excessive compression risks inducing operational ambiguity, necessitating careful calibration.
VII-B Healthcare and Medical Imaging
@@
The effective synthesis of multimodal medical data is pivotal to advancing contemporary medical diagnosis and research. MLLMs can integrate radiographic findings, medical histories, and ancillary diagnostic tests to generate differential diagnoses, which clinicians can correlate with patient records and physician notes to enhance diagnostic accuracy [200]. Furthermore, MLLMs can automatically draft preliminary radiology reports, potentially reducing the workload of radiologists [201, 202, 203]. A major challenge for MLLMs in medical imaging is the processing of high-resolution images, such as Whole-slide Images (WSIs) in pathology, which can contain billions of pixels. To overcome this, TCP-LLaVA [204] uses a set of trainable compression tokens to aggregate and condense crucial information from thousands of visual and textual inputs. Instead of feeding every single image patch token into the language model, only these compressed tokens are forwarded for answer generation. Token compression holds vast potential for widespread application in the field of healthcare and medical imaging, enabling the efficient analysis of complex, high-resolution images.
VII-C Robotics and Autonomous Systems
@@
Leveraging the significant capabilities of video LLMs in long-form video comprehension enables their deployment in robotics [42] and autonomous driving systems [43, 205]. However, the inherent computational complexity of long-duration video processing creates fundamental latency-efficiency tradeoffs that challenge real-time implementation. Token compression addresses this by prioritizing salient spatio-temporal dynamics (e.g., agent movements, action trajectories) and fine-grained per-frame details, enabling computationally efficient video understanding for these domains. VTS [43] proposes a token pruning strategy for autonomous driving scenarios. VTS employs a proposal model based on a lightweight convolutional neural network that is able to adaptively identify keyframes and pry less informative tokens (e.g., invariant backgrounds and stationary objects). StreamVLN [42] further enhances inference efficiency for real-time navigation by employing a voxel-based spatial pruning strategy at test time to reduce memory tokens. This approach makes real-time navigation feasible.
VII-D Efficient Reasoning
@@
Token compression improves efficiency by removing redundant input tokens. However, in many cases, the main source of computational cost shifts from the input to the output, most notably in reasoning models [206, 207, 208], where lengthy generation chains are common. The “slow-thinking” paradigm improves reasoning ability but results in lengthy reasoning chains [209, 210, 211, 212]. Some efficient reasoning methods compress these chains using similar techniques (e.g., attention mechanisms, semantic importance) [213, 214, 215, 216], typically requiring fine-tuning via Supervised Fine-Tuning (SFT) or Reinforcement Learning (RL). Beyond token compression, other approaches improve reasoning efficiency by compressing the model [217, 218, 219, 220] or accelerating decoding [221, 222, 223, 224, 225].
VIII Conclusion
This paper presents the first structured survey of token compression techniques for MLLMs. This is a critical area, driven by the growing need to process extensive and intricate data across various modalities. We systematically categorized existing methods based on their primary data focus: image-centric, video-centric, and audio-centric compression, acknowledging that effective strategies are intrinsically linked to the unique characteristics and redundancies of each modality. Furthermore, we delved into the underlying mechanisms of these methods, including transformation-based, similarity-based, attention-based, and query-based approaches. By consolidating current progress and identifying key challenges, this survey aims to inspire future research and accelerate advancements in this rapidly evolving field. The insights presented are designed to guide researchers in developing more efficient and scalable MLLMs, ultimately pushing the boundaries of what these powerful models can achieve with increasingly longer and more complex inputs.

### v5

@@ 6.5.4 Evaluation Challenges
For training-free token compression methods, current evaluations often rely on metrics such as the number of compressed tokens and FLOPs. However, these metrics offer an incomplete picture. While the number of compressed tokens provides a preliminary classification, the compression location significantly impacts the downstream computational load; earlier pruning generally leads to greater reductions. Similarly, FLOPs, while useful for theoretical computational estimates, frequently do not accurately reflect actual inference speed. Therefore, for training-free methods, more practical metrics like Time To First Token (TTFT) and decoding latency per token are crucial for a more accurate assessment of real-world inference acceleration.
Evaluation Benchmarks Gap. Current evaluation datasets for MLLM token compression often rely on general multimodal benchmarks (Table 4), provide insufficient granularity. For example, in challenging long video understanding tasks, performance hinges more on sparse frame sampling, capturing key frames, than on the specific token compression method. This can obscure the true impact of token compression, making its efficacy appear negligible. Furthermore, relying solely on VQA datasets that demand only low-fidelity information is insufficient, as they lack the fine-grained sensitivity required to evaluate token compression.
This reveals a critical gap: current datasets often fail to isolate and precisely measure the effect of token compression. Therefore, adopting specific designed evaluation methodologies like EffiVLM-Bench (Wang et al., 2025c), VTC-Bench (Liao et al., 2025) and challenging benchmarks such as OCR (Yu et al., 2016) and temporal grounding (Gao et al., 2017) benchmarks, is crucial for accurately assessing the true efficacy and nuanced benefits of token compression methods.
6.6 Pruning Location and Trade-offs
@@
6.7 Future Directions
6.7.1 Joint Token Compression for Multimodal Settings
While distinct modalities exhibit unique redundancy patterns requiring specialized handling, the field is rapidly evolving towards Omnimodal Large Language Models (omni LLMs) capable of real-time, joint inference (Xu et al., 2025b; Tong et al., 2025; Xu et al., 2025c; Xie and Wu, 2024; Tang et al., 2025; Yang et al., 2025b; Ge et al., 2025; Fu et al., 2024; Shu et al., 2025; Sun et al., 2024a; Li et al., 2024b). However, single-modal deployments remain constrained by their unimodal inputs. As established in Sections 3, 4, and 5, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. A promising future direction lies in exploiting cross-modal synergy to reduce the aggregate token count. Pioneering efforts like OmniZip (Tao et al., 2025b) have begun to explore this by utilizing audio cues to guide visual pruning, underscoring the predictive utility of one modality over another. Future research should further investigate deep joint compression mechanisms, where the redundancy of audio, video, and textual tokens is evaluated holistically, enabling efficient, long-context interaction for next-generation omni LLMs.
6.7.2 Improved Architecture
Current token compression methods are often employed as a remedial measure to process long contexts efficiently. However, a more valuable approach might involve designing model architectures that intrinsically account for data redundancy during their initial conception. By doing so, the number of tokens could be reduced during the abstraction of data features. This is particularly relevant for current architectures, especially those of video LLMs, where generated tokens still exhibit significant redundancy. Therefore, exploring architectural designs that inherently foster more condensed information abstraction from the outset represents a promising research direction.
Furthermore, recent architectures utilizing linear attention (Gu and Dao, 2024b; Peng et al., 2023; Sun et al., 2023; Qiu et al., 2025) have emerged as a parallel solution, mitigating the computational explosion associated with increasing token counts through linear complexity. However, determining how to effectively identify and eliminate input redundancy within these novel frameworks to achieve more compact data representations remains a promising avenue for future exploration.
7 Applications
@@
7.1 GUI Agents and Human-Computer Interaction
@@
Graphical user interface (GUI) agents perceive and interact with visual interfaces, interpret natural language instructions, analyze GUI states, and execute corresponding actions. These agents have to parse screen streams in real-time, producing extensive token sequences that often exceed computational limits (Zhang et al., 2024b; Wang et al., 2024a). Multimodal token compression enhances the efficiency of GUI agents. This approach mitigates context overflow in extended operation sequences by dynamically compressing redundant visual elements (e.g., extra white space or simple backgrounds). For some small but important control elements, it should also eliminate other irrelevant visual elements and highlight their importance. For instance, ShowUI (Lin et al., 2025a) is the first model to apply token selection strategy to GUI agents. ShowUI segments GUI screenshots into connected components by clustering pixels with similar RGB values, significantly reducing the total number of discrete elements. During both training and inference phases, the system employs an adaptive token selection strategy that probabilistically prunes redundant tokens within these components, thereby optimizing computational efficiency while preserving functional semantics However, excessive compression risks inducing operational ambiguity, necessitating careful calibration.
7.2 Healthcare and Medical Imaging
@@
The effective synthesis of multimodal medical data is pivotal to advancing contemporary medical diagnosis and research. MLLMs can integrate radiographic findings, medical histories, and ancillary diagnostic tests to generate differential diagnoses, which clinicians can correlate with patient records and physician notes to enhance diagnostic accuracy (Liang et al., 2024). Furthermore, MLLMs can automatically draft preliminary radiology reports, potentially reducing the workload of radiologists (Beddiar and Oussalah, 2023; Bazi et al., 2023; He et al., 2020). A major challenge for MLLMs in medical imaging is the processing of high-resolution images, such as Whole-slide Images (WSIs) in pathology, which can contain billions of pixels. To overcome this, TCP-LLaVA (Lyu et al., 2025) uses a set of trainable compression tokens to aggregate and condense crucial information from thousands of visual and textual inputs. Instead of feeding every single image patch token into the language model, only these compressed tokens are forwarded for answer generation. Token compression holds vast potential for widespread application in the field of healthcare and medical imaging, enabling the efficient analysis of complex, high-resolution images.
7.3 Robotics and Autonomous Systems
@@
Leveraging the significant capabilities of video LLMs in long-form video comprehension enables their deployment in robotics (Wei et al., 2025) and autonomous driving systems (Ma et al., 2024b; Zhou et al., 2024; Zhu et al., 2025b). However, the inherent computational complexity of long-duration video processing creates fundamental latency-efficiency tradeoffs that challenge real-time implementation. Token compression addresses this by prioritizing salient spatio-temporal dynamics (e.g., agent movements, action trajectories) and fine-grained per-frame details, enabling computationally efficient video understanding for these domains. VTS (Ma et al., 2024b) proposes a token pruning strategy for autonomous driving scenarios. VTS employs a proposal model based on a lightweight convolutional neural network that is able to adaptively identify keyframes and pry less informative tokens (e.g., invariant backgrounds and stationary objects). StreamVLN (Wei et al., 2025) further enhances inference efficiency for real-time navigation by employing a voxel-based spatial pruning strategy at test time to reduce memory tokens. This approach makes real-time navigation feasible.
7.4 Efficient Reasoning
@@
Token compression improves efficiency by removing redundant input tokens. However, in many cases, the main source of computational cost shifts from input to output, most notably in reasoning models (Team et al., 2025; Guo et al., 2025; Jaech et al., 2024), where lengthy generation chains are common. The “slow-thinking” paradigm improves reasoning ability but results in lengthy reasoning chains (Feng et al., 2025a; Sui et al., 2025; Chen et al., 2025; Feng et al., 2025c; b). Some efficient reasoning methods compress these chains using similar techniques (e.g., attention mechanisms, semantic importance) (Ma et al., 2025a; Xia et al., 2025; Liu et al., 2024b; Fang et al., 2025), typically requiring fine-tuning via Supervised Fine-Tuning (SFT) or Reinforcement Learning (RL). Beyond token compression, other approaches improve reasoning efficiency by compressing model (Magister et al., 2022; Li et al., 2023b; Feng et al., 2024; Zhang et al., 2025d) or accelerating decoding (Sun et al., 2024c; Ma et al., 2024a; Luo et al., 2025; Xu et al., 2025a; Ding et al., 2025).
8 Conclusion
This paper presents the first structured survey of token compression techniques for Multimodal Large Language Models (MLLMs), establishing a taxonomy based on modality-specific redundancy and underlying compression mechanisms. While current methods demonstrate promising efficiency gains, several critical challenges remain on the path toward scalable and robust MLLMs. Future research must move beyond simple redundancy reduction to address the preservation of cross-modal alignment under high compression ratios and the maintenance of causal reasoning capabilities in temporal sequences. Furthermore, the field necessitates the development of specialized benchmarks designed to rigorously evaluate multi-frame comprehension and long-term context retention. We hope this survey serves as a roadmap, guiding the community to tackle these open problems and push the boundaries of processing increasingly complex multimodal data.
9 Acknowledgment
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of National Talent Program and Hangzhou Municipal Talent Program.